OpenStudy (anonymous):

A company distributes free candies to all the students of x schools. Each school has (x + 1) classes. The number of students in each class is 3 more than the number of classes in each school. Each student is given 4 candies. Part A: Write an expression to show the total number of candies distributed by the company in x schools. (4 points) Part B: What would x(x + 1) represent? When simplified, what would be the degree and classification of this expression? (4 points) Part C: How can you calculate the total number of students in each school? (2 points)

OpenStudy (anonymous):

Part A:So the total number of candies is given 4 each to x+4 students in x+1 classes in x schools. total number of candies = 4(x+4)(x+1)x Part B: x is the number schools x+1 is the number of classes in each school x(x+1) would represent the total number of classes all taken together. When simplified x(x+1) = x^2+1 which is a second degree polynomial.

OpenStudy (anonymous):

i need help on part C @perl

OpenStudy (anonymous):

@amistre64 @myininaya

OpenStudy (anonymous):

@TuringTest @Michele_Laino

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (studygurl14):

To calculate the total number of students, you would use the expression without the candy part. So x(x+1)(x+4)

OpenStudy (studygurl14):

BTW, nice work with Parts A and B